A sample of 25 elements from a normally distributed population is selected.the sample mean is 10.assuming a population standard deviationog 4, determine the 95%confidence interval for mean ?


1
Expert's answer
2022-03-03T13:58:20-0500

"95\%CI=(\bar x-z_{0.025}\frac{s}{\sqrt{n}},\bar x+z_{0.025}\frac{s}{\sqrt{n}})="

"=(10-1.96\frac{4}{\sqrt{25}},10+1.96\frac{4}{\sqrt{25}})=(8.432,11.568)."